2 times a number, increased by 28 is less than or eaqual to 6 times that number.
Answer by tutorcecilia(2152)

Let x= a number
2x+28<=6x
2x-2x+28<=6x-2x
28<=4x
28/4<=4x/4
7<=x
or x=>7